### v2

@@ 4.1 Implications of the Aubin property
Note that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Hence, (4.14) holds if and only if for all ${\bm{\nu}}\in{\mathds{R}}^{\ell}\setminus\{{\boldsymbol{0}}\}$ it holds that
$\begin{array}[]{ll}\Theta H{\bm{\nu}}\notin{\mathcal{R}}-\left\{\sum\limits_{j\in J_{4,5,6}}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}\ \Big{|}\ {\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}})\right\}.\end{array}$ (4.15)
Therefore, the proof of the proposition is finished if (4.13) and (4.15) are equivalent.
On the one hand, suppose that (4.13) does not hold, i.e., there exists a nonzero vector $\tilde{\bm{\nu}}\in{\mathds{R}}^{\ell}$ such that
$H^{\top}\Theta H\tilde{\bm{\nu}}=-H^{\top}\sum_{j\in J_{4,5,6}}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}\ \mbox{with}\ {\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H\tilde{\bm{\nu}}).$ (4.16)
Therefore, one has $\Theta H\tilde{\bm{\nu}}+\sum_{j\in J_{4,5,6}}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}\in\ker H^{\top}$. Let $\Xi$ be the matrix defined in (4.10). Then, for any ${\bm{\zeta}}^{0}\in{\mathds{R}}^{m}$, $\{{\bm{\zeta}}^{j}\in{\mathds{R}}^{1+r_{j}}\}_{j\in J_{1}}$, $\{\zeta^{j}\in{\mathds{R}}\}_{j\in J_{3}}$, and $\{\hat{\bm{\zeta}}^{j}\in{\mathds{R}}^{r_{j}}\}_{j\in J_{4}}$ one has
@@
Since $\ker(H^{\top})={{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))}^{\perp}={\sf rge}\Xi^{\top}$, by defining ¯$\overline{\mathcal{R}}:={\mathcal{R}}\cup{\sf span}(\{\nabla g^{j}({\boldsymbol{x}}^{*})S_{j}\}_{j\in J_{4}})$ one can get
@@
Note that for any $j\in J_{4}$ one has from (4.1) that $({\bm{\mu}}^{*})^{j}\in{\sf bdry}{\mathcal{Q}}_{j}\setminus\{{\boldsymbol{0}}\}$. Meanwhile, as $H\tilde{\bm{\nu}}\in{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$, one has from (4.4) that
@@
Recall from (4.16) that ${\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H\tilde{\bm{\nu}})$. Meanwhile, for any $j\in J_{4}$, it holds that $\langle{\boldsymbol{p}}^{j}-S_{j}\hat{\bm{\zeta}}^{j},{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H\tilde{\bm{\nu}}\rangle=\langle{\boldsymbol{p}}^{j},{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H\tilde{\bm{\nu}}\rangle$. Then, it is easy to see from Part (5) of Lemma 3.2 that
@@
Consequently, (4.15) does not hold.
On the other hand, suppose that (4.15) is not true. Then there are vectors and real numbers $\hat{\bm{\nu}}\in{\mathds{R}}^{\ell}$, ${\bm{\zeta}}^{0}\in{\mathds{R}}^{m}$, $\{{\bm{\zeta}}^{j}\in{\mathds{R}}^{1+r_{j}}\}_{j\in J_{1}}$, $\{\zeta^{j}\in{\mathds{R}}\}_{j\in J_{3}}$, and $\{{\boldsymbol{p}}^{j}\in{\mathds{R}}^{1+r_{j}}\}_{j\in J_{4,5,6}}$ such that
@@
with ${\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H\hat{\bm{\nu}})$ for all $j\in J_{4,5,6}$. Then, for any ${\bm{\nu}}\in{\mathds{R}}^{\ell}$, by noting that $H{\bm{\nu}}\in{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$ and using (4.4) one can get
@@
Since ${\bm{\nu}}$ can be arbitrarily chosen, one can get $H^{\top}\Theta H\hat{\bm{\nu}}=-\sum_{j\in J_{4,5,6}}H^{\top}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}$, which contradicts (4.13). This completes the proof. ∎
@@
Suppose that Assumption 4.1 holds. Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix given in (4.12), and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Then the mapping ${{\mathbb{S}}}_{\rm GE}$ in (1.8) having the Aubin property at the origin for ${\boldsymbol{x}}^{*}$ implies that for all ${\bm{\nu}}\in{\mathds{R}}^{\ell}\setminus\{{\boldsymbol{0}}\}$ it holds that
$\begin{array}[]{ll}H^{\top}\Theta H{\bm{\nu}}\notin-\left\{\sum\limits_{j\in J_{4,5}}{\mathcal{G}}_{j}^{\top}p^{j}+\sum\limits_{j\in J_{6}}{\mathcal{G}}_{j}^{\top}{\boldsymbol{p}}^{j}\ \Big{|}\ \begin{array}[]{ll}p^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j}{\bm{\nu}}),\\[2.84526pt] {\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j}{\bm{\nu}})\end{array}\right\},\end{array}$ (4.17)
@@
${\mathcal{G}}_{j}:=\begin{cases}((\dot{\mu}^{*})^{j};-(\bar{\bm{\mu}}^{*})^{j})^{\top}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H,&\forall\,j\in J_{4},\\ (\dot{g}^{j}({\boldsymbol{x}}^{*});-\bar{g}^{j}({\boldsymbol{x}}^{*}))^{\top}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H,&\forall\,j\in J_{5},\\ {\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H,&\forall\,j\in J_{6}.\end{cases}$ (4.18)
@@
Proof.
Let ${\boldsymbol{0}}\neq{\bm{\nu}}\in{\mathds{R}}^{\ell}$ be arbitrarily chosen. For $j\in J_{4}$, one has from (4.4) that ${\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}}=\tau_{j}((\dot{\mu}^{*})^{j};-(\bar{\bm{\mu}}^{*})^{j})$ with $\tau_{j}\in{\mathds{R}}$. Meanwhile, from (4.1) one has $g^{j}({\boldsymbol{x}}^{*})={\boldsymbol{0}}$ and $({\bm{\mu}}^{*})^{j}\in{\sf bdry}{\mathcal{Q}}_{j}\setminus\{{\boldsymbol{0}}\}$. Then by Part (5) of Lemma 3.2 and (4.18) one has ${\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}})$ if and only if
@@
which, by (2.4), is equivalent to $-\langle{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}},{\boldsymbol{p}}^{j}\rangle\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j}{\bm{\nu}})$.
For $j\in J_{5}$, from (4.1) one has $g^{j}({\boldsymbol{x}}^{*})\in{\sf bdry}{\mathcal{Q}}_{j}\setminus\{{\boldsymbol{0}}\}$ and $({\bm{\mu}}^{*})^{j}={\boldsymbol{0}}$. Therefore, by Part (4) of Lemma 3.2 one has ${\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}})$ if and only if ${\boldsymbol{p}}^{j}=p^{j}(\dot{g}^{j}({\boldsymbol{x}}^{*});-\bar{g}^{j}({\boldsymbol{x}}^{*}))$ with
@@
By (2.4), the above condition on $p^{j}$ is exactly $p^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j}{\bm{\nu}})$. Furthermore, one has $g^{j}({\boldsymbol{x}}^{*})=({\bm{\mu}}^{*})^{j}={\boldsymbol{0}}$ for $j\in J_{6}$. Therefore, (4.13) in Lemma 4.1 is equivalent to
$\begin{array}[]{ll}H^{\top}\Theta H{\bm{\nu}}\notin-\left\{\begin{array}[]{ll}\sum\limits_{j\in J_{4}}H^{\top}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}+\sum\limits_{j\in J_{5}}{\mathcal{G}}_{j}^{\top}p^{j}+\sum\limits_{j\in J_{6}}{\mathcal{G}}_{j}^{\top}{\boldsymbol{p}}^{j}\ \Big{|}\\[8.53581pt] \hfill-\langle{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}},{\boldsymbol{p}}^{j}\rangle\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j}{\bm{\nu}}),\ j\in J_{4},\\[5.69054pt] \hfill p^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j}{\bm{\nu}}),\ j\in J_{5},\\[5.69054pt] \hfill{\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j}{\bm{\nu}}),\ j\in J_{6}\end{array}\right\}.\end{array}$ (4.19)
@@
Therefore, (4.17) is a consequence of (4.19).
Suppose that there exists a nonzero vector ${\bm{\nu}}\in{\mathds{R}}^{l}$ such that $H^{\top}\Theta H{\bm{\nu}}={\boldsymbol{0}}$. In this case, one can take ${\boldsymbol{p}}^{j}={\boldsymbol{0}}$ for all $j\in J_{4,6}$ and $p^{j}=0$ for all $j\in J_{5}$ to get valid coderivatives in (4.19) via (2.4) and Part (6e) of Lemma 3.2. This contradicts (4.19), so the matrix $H^{\top}\Theta H$ is nonsingular, which completes the proof. ∎
@@
Suppose that Assumption 4.1 holds. Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix defined in (4.12), and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. For ${\mathcal{G}}_{j}$ defined in (4.18), one has
${\sf rge}\left(\{{\mathcal{G}}_{j}\}_{{j\in J_{4,5,6}}}\right)={\mathds{R}}^{|J_{4,5}|+\sum_{j\in J_{6}}(1+r_{j})},$ (4.20)
@@
Proof.
For the given vectors $\zeta^{j}\in{\mathds{R}}$ ($j\in J_{4,5}$) and ${\bm{\zeta}}^{j}\in{\mathds{R}}^{1+r_{j}}$ ($j\in J_{6}$), we show that there exists a vector ${\bm{\nu}}\in{\mathds{R}}^{\ell}$ such that ${\mathcal{G}}_{j}{\bm{\nu}}=\zeta^{j}$ for all $j\in J_{4,5}$ and ${\mathcal{G}}_{j}{\bm{\nu}}={\bm{\zeta}}^{j}$ for all $j\in J_{6}$. Recall that the constraint nondegeneracy condition (1.7) holds at ${\boldsymbol{x}}^{*}$. Then, for the given vectors ${\boldsymbol{z}}^{j}\in{\mathds{R}}^{1+r_{j}}$ defined by
@@
where ${\sf lin}\left({\mathcal{T}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}))\right)$ is defined in (4.11). It can be routinely examined from (4.4) that ${\boldsymbol{d}}\in{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Moreover, since $H$ has full column rank such that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$, there exists a unique vector ${\bm{\nu}}\in{\mathds{R}}^{\ell}$ such that $H{\bm{\nu}}={\boldsymbol{d}}$. For all $j\in J_{4}$, one has ${\boldsymbol{v}}^{j}={\boldsymbol{0}}$, so that by (4.18) one can get ${\mathcal{G}}_{j}{\bm{\nu}}=((\dot{\mu}^{*})^{j};-(\bar{\bm{\mu}}^{*})^{j})^{\top}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}}=\zeta^{j}$. Furthermore, for $j\in J_{5}$, one has $\langle\bar{\boldsymbol{v}}^{j},\bar{g}^{j}({\boldsymbol{x}}^{*})\rangle-\dot{v}^{j}\dot{g}^{j}({\boldsymbol{x}}^{*})=0$. Thus by (4.18) one has
@@
Finally, for $j\in J_{6}$, one has from (4.18) that ${\boldsymbol{v}}^{j}={\boldsymbol{0}}$ and ${\mathcal{G}}_{j}{\bm{\nu}}={\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}}={\bm{\zeta}}^{j}$. Consequently, we know that (4.20) holds, which completes the proof. ∎
Note that in the reduction procedure introduced above, we have reduced the original condition (4.9) for characterizing the Aubin property of ${\mathbb{S}}_{\rm GE}$ to the condition (4.17) in Corollary 4.1. The former condition is related to all the constraints in the nonlinear SOCP (1.1), while the latter condition is only related to the blocks at which the strict complementarity does not hold.

### v3

@@ 4.1 Implications of the Aubin property
Note that ${\sf rge}H={\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Hence, (40) holds if and only if for all ${\bm{\nu}}\in{\mathds{R}}^{\ell}\setminus\{{\boldsymbol{0}}\}$ it holds that
(41) $\begin{array}[]{ll}\Theta H{\bm{\nu}}\notin{\mathcal{R}}-\left\{\sum\limits_{j\in J_{4,5,6}}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}\ \Big{|}\ {\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}})\right\}.\end{array}$
Therefore, the proof of the proposition is finished if (39) and (41) are equivalent.
On the one hand, suppose that (39) does not hold, i.e., there exists a nonzero vector $\tilde{\bm{\nu}}\in{\mathds{R}}^{\ell}$ such that
(42) $H^{\top}\Theta H\tilde{\bm{\nu}}=-H^{\top}\sum_{j\in J_{4,5,6}}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}\ \mbox{with}\ {\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H\tilde{\bm{\nu}}).$
Therefore, one has $\Theta H\tilde{\bm{\nu}}+\sum_{j\in J_{4,5,6}}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}\in{\sf ker}H^{\top}$. Let $\Xi$ be the matrix defined in (36). Then, for any ${\bm{\zeta}}^{0}\in{\mathds{R}}^{m}$, $\{{\bm{\zeta}}^{j}\in{\mathds{R}}^{1+r_{j}}\}_{j\in J_{1}}$, $\{\zeta^{j}\in{\mathds{R}}\}_{j\in J_{3}}$, and $\{\hat{\bm{\zeta}}^{j}\in{\mathds{R}}^{r_{j}}\}_{j\in J_{4}}$ one has
@@
Since ${\sf ker}(H^{\top})={{\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))}^{\perp}={\sf rge}\Xi^{\top}$, by defining ¯$\overline{\mathcal{R}}:={\mathcal{R}}\cup{\sf span}(\{\nabla g^{j}({\boldsymbol{x}}^{*})S_{j}\}_{j\in J_{4}})$ one can get
@@
Note that for any $j\in J_{4}$ one has from (27) that $({\bm{\mu}}^{*})^{j}\in{\sf bdry}{\mathcal{Q}}_{j}\setminus\{{\boldsymbol{0}}\}$. Meanwhile, as $H\tilde{\bm{\nu}}\in{\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$, one has from (30) that
@@
Recall from (42) that ${\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H\tilde{\bm{\nu}})$. Meanwhile, for any $j\in J_{4}$, it holds that $\langle{\boldsymbol{p}}^{j}-S_{j}\hat{\bm{\zeta}}^{j},{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H\tilde{\bm{\nu}}\rangle=\langle{\boldsymbol{p}}^{j},{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H\tilde{\bm{\nu}}\rangle$. Then, it is easy to see from Part (5) of Lemma 3.2 that
@@
Consequently, (41) does not hold.
On the other hand, suppose that (41) is not true. Then there are vectors and real numbers $\hat{\bm{\nu}}\in{\mathds{R}}^{\ell}$, ${\bm{\zeta}}^{0}\in{\mathds{R}}^{m}$, $\{{\bm{\zeta}}^{j}\in{\mathds{R}}^{1+r_{j}}\}_{j\in J_{1}}$, $\{\zeta^{j}\in{\mathds{R}}\}_{j\in J_{3}}$, and $\{{\boldsymbol{p}}^{j}\in{\mathds{R}}^{1+r_{j}}\}_{j\in J_{4,5,6}}$ such that
@@
with ${\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H\hat{\bm{\nu}})$ for all $j\in J_{4,5,6}$. Then, for any ${\bm{\nu}}\in{\mathds{R}}^{\ell}$, by noting that $H{\bm{\nu}}\in{\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$ and using (30) one can get
@@
Since ${\bm{\nu}}$ can be arbitrarily chosen, one can get $H^{\top}\Theta H\hat{\bm{\nu}}=-\sum_{j\in J_{4,5,6}}H^{\top}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}$, which contradicts (39). This completes the proof.
@@
Suppose that Assumption 4.1 holds. Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix given in (38), and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Then the mapping ${{\mathbb{S}}}_{\rm GE}$ in (8) having the Aubin property at the origin for ${\boldsymbol{x}}^{*}$ implies that for all ${\bm{\nu}}\in{\mathds{R}}^{\ell}\setminus\{{\boldsymbol{0}}\}$ it holds that
(43) $\begin{array}[]{ll}H^{\top}\Theta H{\bm{\nu}}\notin-\left\{\sum\limits_{j\in J_{4,5}}{\mathcal{G}}_{j}^{\top}p^{j}+\sum\limits_{j\in J_{6}}{\mathcal{G}}_{j}^{\top}{\boldsymbol{p}}^{j}\ \Big{|}\ \begin{array}[]{ll}p^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j}{\bm{\nu}}),\\[2.84526pt] {\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j}{\bm{\nu}})\end{array}\right\},\end{array}$
@@
(44) ${\mathcal{G}}_{j}:=\begin{cases}((\dot{\mu}^{*})^{j};-(\bar{\bm{\mu}}^{*})^{j})^{\top}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H,&\forall\,j\in J_{4},\\ (\dot{g}^{j}({\boldsymbol{x}}^{*});-\bar{g}^{j}({\boldsymbol{x}}^{*}))^{\top}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H,&\forall\,j\in J_{5},\\ {\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H,&\forall\,j\in J_{6}.\end{cases}$
@@
Proof 4.10.
Let ${\boldsymbol{0}}\neq{\bm{\nu}}\in{\mathds{R}}^{\ell}$ be arbitrarily chosen. For $j\in J_{4}$, one has from (30) that ${\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}}=\tau_{j}((\dot{\mu}^{*})^{j};-(\bar{\bm{\mu}}^{*})^{j})$ with $\tau_{j}\in{\mathds{R}}$. Meanwhile, from (27) one has $g^{j}({\boldsymbol{x}}^{*})={\boldsymbol{0}}$ and $({\bm{\mu}}^{*})^{j}\in{\sf bdry}{\mathcal{Q}}_{j}\setminus\{{\boldsymbol{0}}\}$. Then by Part (5) of Lemma 3.2 and (44) one has ${\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}})$ if and only if
@@
which, by (12), is equivalent to $-\langle{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}},{\boldsymbol{p}}^{j}\rangle\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j}{\bm{\nu}})$.
For $j\in J_{5}$, from (27) one has $g^{j}({\boldsymbol{x}}^{*})\in{\sf bdry}{\mathcal{Q}}_{j}\setminus\{{\boldsymbol{0}}\}$ and $({\bm{\mu}}^{*})^{j}={\boldsymbol{0}}$. Therefore, by Part (4) of Lemma 3.2 one has ${\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}})$ if and only if ${\boldsymbol{p}}^{j}=p^{j}(\dot{g}^{j}({\boldsymbol{x}}^{*});-\bar{g}^{j}({\boldsymbol{x}}^{*}))$ with
@@
By (12), the above condition on $p^{j}$ is exactly $p^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j}{\bm{\nu}})$. Furthermore, one has $g^{j}({\boldsymbol{x}}^{*})=({\bm{\mu}}^{*})^{j}={\boldsymbol{0}}$ for $j\in J_{6}$. Therefore, (39) in Lemma 4.7 is equivalent to
(45) $\begin{array}[]{ll}H^{\top}\Theta H{\bm{\nu}}\notin-\left\{\begin{array}[]{ll}\sum\limits_{j\in J_{4}}H^{\top}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}+\sum\limits_{j\in J_{5}}{\mathcal{G}}_{j}^{\top}p^{j}+\sum\limits_{j\in J_{6}}{\mathcal{G}}_{j}^{\top}{\boldsymbol{p}}^{j}\ \Big{|}\\[8.53581pt] \hfill-\langle{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}},{\boldsymbol{p}}^{j}\rangle\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j}{\bm{\nu}}),\ j\in J_{4},\\[5.69054pt] \hfill p^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{[0,+\infty)}(0,0)({\mathcal{G}}_{j}{\bm{\nu}}),\ j\in J_{5},\\[5.69054pt] \hfill{\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}({\boldsymbol{0}},{\boldsymbol{0}})({\mathcal{G}}_{j}{\bm{\nu}}),\ j\in J_{6}\end{array}\right\}.\end{array}$
@@
Therefore, (43) is a consequence of (45).
Suppose that there exists a nonzero vector ${\bm{\nu}}\in{\mathds{R}}^{l}$ such that $H^{\top}\Theta H{\bm{\nu}}={\boldsymbol{0}}$. In this case, one can take ${\boldsymbol{p}}^{j}={\boldsymbol{0}}$ for all $j\in J_{4,6}$ and $p^{j}=0$ for all $j\in J_{5}$ to get valid coderivatives in (45) via (12) and Part (6e) of Lemma 3.2. This contradicts (45), so the matrix $H^{\top}\Theta H$ is nonsingular, which completes the proof.
@@
Suppose that Assumption 4.1 holds. Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix defined in (38), and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. For ${\mathcal{G}}_{j}$ defined in (44), one has
(46) ${\sf rge}\left(\{{\mathcal{G}}_{j}\}_{{j\in J_{4,5,6}}}\right)={\mathds{R}}^{|J_{4,5}|+\sum_{j\in J_{6}}(1+r_{j})},$
@@
Proof 4.12.
For the given vectors $\zeta^{j}\in{\mathds{R}}$ ($j\in J_{4,5}$) and ${\bm{\zeta}}^{j}\in{\mathds{R}}^{1+r_{j}}$ ($j\in J_{6}$), we show that there exists a vector ${\bm{\nu}}\in{\mathds{R}}^{\ell}$ such that ${\mathcal{G}}_{j}{\bm{\nu}}=\zeta^{j}$ for all $j\in J_{4,5}$ and ${\mathcal{G}}_{j}{\bm{\nu}}={\bm{\zeta}}^{j}$ for all $j\in J_{6}$. Recall that the constraint nondegeneracy condition (7) holds at ${\boldsymbol{x}}^{*}$. Then, for the given vectors ${\boldsymbol{z}}^{j}\in{\mathds{R}}^{1+r_{j}}$ defined by
@@
where ${\sf lin}\left({\mathcal{T}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}))\right)$ is defined in (37). It can be routinely examined from (30) that ${\boldsymbol{d}}\in{\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Moreover, since $H$ has full column rank such that ${\sf rge}H={\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$, there exists a unique vector ${\bm{\nu}}\in{\mathds{R}}^{\ell}$ such that $H{\bm{\nu}}={\boldsymbol{d}}$. For all $j\in J_{4}$, one has ${\boldsymbol{v}}^{j}={\boldsymbol{0}}$, so that by (44) one can get ${\mathcal{G}}_{j}{\bm{\nu}}=((\dot{\mu}^{*})^{j};-(\bar{\bm{\mu}}^{*})^{j})^{\top}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}}=\zeta^{j}$. Furthermore, for $j\in J_{5}$, one has $\langle\bar{\boldsymbol{v}}^{j},\bar{g}^{j}({\boldsymbol{x}}^{*})\rangle-\dot{v}^{j}\dot{g}^{j}({\boldsymbol{x}}^{*})=0$. Thus by (44) one has
@@
Finally, for $j\in J_{6}$, one has from (44) that ${\boldsymbol{v}}^{j}={\boldsymbol{0}}$ and ${\mathcal{G}}_{j}{\bm{\nu}}={\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}}={\bm{\zeta}}^{j}$. Consequently, we know that (46) holds, which completes the proof.
Note that in the reduction procedure introduced above, we have reduced the original condition (35) for characterizing the Aubin property of ${\mathbb{S}}_{\rm GE}$ to the condition (43) in Corollary 4.9. The former condition is related to all the constraints in the nonlinear SOCP (1), while the latter condition is only related to the blocks at which the strict complementarity does not hold.